A 13-year, 6 percent coupon bond pays interest semiannually. The bond has a face value of $1,000. What is the percentage change in the price of this bond if the market yield to maturity rises to 5.7 percent from the

current rate of 5.5 percent?

A) 1.97 percent increaseB) 1.79 percent increaseC) 1.79 percent decreaseD) 1.6 percent decreaseE) 1 percent decrease

I need to solve it with this calcualter , please, ( Texas Instruments - BA II Plus) step by step.

Thank you so much for your time and your help
Answer #1

First find price at 5.5%

Assign the following values

FV -1000

PMT -30

PV 0

I/Y 5.5/2 = 2.75%

N 26

CPT PV = 1046

Now find bond price at 5.7%

Just change I/Y to 5.7/2 = 2.85%

CPT PV = 1027.28

CHANGE in price = (1027.28-1046)/1046 = 1.79% decrease

Option C is the answer